Solve the following LPP by the two-phase simplex method:
Maximise Z=x_1+2x_2+3x_3
Subject to
x_1+2x_2+3x_3=15
2x_1+x_2+5x_3=20
x_1+2x_2+x_3+x_4=10
x_1,x_2,x_3,x_4≥0
1
Expert's answer
2019-03-08T08:14:02-0500

1) Introduce slack variables "x_5, x_6, x_7," for the first, the second, and the third equation. then write the function in the form

"Z = - Mx_5 - Mx_6 - Mx_7 \u2192 max ,"

express the slack variables:


"x_5 = 15-x_1-2x_2-3x_3""x_6 = 20-2x_1-x_2-5x_3""x_7 = 10-x_1-2x_2-x_3-x_4"

Substitute them to the function:


"Z= 4Mx_1+5Mx_2+9Mx_3+Mx_4-45M\u2192 max ,"

After the first phase (common simplex-method with optimal solution) has ended at the following matrix:


"\begin{matrix}
 0 & -1 & 0 & 0& 0 \\
 15/7 & -1/7 & 0 & 0& 0 \\
 25/7 & 3/7& 0 & 0 & 0 \\
15/7 & 6/7 & 0 & 0 & 0\\
0 & 0 & -1 & 0 &0 \\
0 & 0 & 0 & -1 & 0 \\
0 & 0 & 0 & 0 & -1 \\
0 & 0 & 0 & 0 & 0 \\

\end{matrix}"

Express the basic variables:


"x_2 = 15/7+1/7x_1, \\
x_3 = 25/7-3/7x_1."

Substitute them to the function:


"Z=x_1 + 2(15\/7+1\/7x_1) + 3(25\/7-3\/7x_1)=15."

Finally we get the correct matrix after removing the rows with slack variables (Phase II):


"\begin{matrix}
 0 & -1 & 0 & 0& 0 \\
 15/7 & -1/7 & 0 & 0& 0 \\
 25/7 & 3/7& 0 & 0 & 0 \\
 15/7 & 6/7 & 0 & 0 & 0\\
15 & 0 & 0 & 0 &0
\end{matrix}"


The two-phase simplex method gives:


"x_1 = 0 \\
x_2 = 15/7 \\
x_3 = 25/7 \\
Z = 2•15/7 + 3•25/7 + 0•15/7 = 15."